OpenStudy (mayaal):

giving off a medal! Verify the identity: cot(x-pi/2)=-tan x

OpenStudy (mathmate):

Hint: \(\huge cot(A-B)=-\frac{1+cot(A)cot(B)}{cot(A)-cot(B)}\)

OpenStudy (abmon98):

\[Cot(A-B)=Cos(A-B)/Sin(A-B) \]\[=CosACosB+SinASinB/SinACosB-sinBCosA \] \[\cos(\pi/2)=0\] \[\sin(\pi/2)=1\]

OpenStudy (mayaal):

Which identity is this? btw,thnks 4 replying:)

OpenStudy (mayaal):

is this a sum and difference formula,@Abmon98 ?

OpenStudy (abmon98):

\[Cotx=1/tanx\] \[tanx=sinx/cosx\] Addition formula thats subtracting two angles at my level i do not know how to proof it i am sorry. i just learn them. Sin(A-B)=SinACosB-SinBCosA Sin(A+B)=SinACosB+SinBCosA Cos(A-B)=CosACosB+SinASinB Cos(A+B)=CosACosB-SinASinB

OpenStudy (abmon98):

\[Cosx*Cos(\pi/2)+Sinx*Sin(\pi/2)/Sinx*Cos(\pi/2)-Sin(\pi/2)*Cosx\] \[-Sinx/cosx=-tanx\]

OpenStudy (mayaal):

It's Ok,thanks for trying to help me:)

OpenStudy (abmon98):

your most welcome :)

OpenStudy (mathmate):

The cot(A-B) identity may not be very well-known, but it can be as powerful: Let B=\(\pi/2\), then cotB=0 \(\large cot(A-B)=-\frac{1+cotA*0}{cotA-0}\) \(\large =-\frac1{cotA}\) =-tanA